Determine the magnitude of the electric field at a point midway between a -8.0μC and a 6.1 μC charge 8.0cm apart. Assume no other charges are nearby.

Express your answer using two significant figures.
Answer #1


q1 = -8*10^-6 C


q2 = 6.1*10^-6 C


d = 8 cm = 0.08 m

x1 = 4 cm = 0.05 m


x2 = 4 cm = 0.04 m


E1 = k*q1/x1^2 directed toward q1

E1 = 9*10^9*8*10^-6/0.04^2 = 45*10^6 N/C


E2 = k*q2/x2^2 directed toward q1

E2 = 9*10^9*6.1*10^-6/0.04^2 = 34*10^6 N/C

total eelctric field E E1 + E2 = 79*10^6 N/C <<<<----------answer